A rectangle has twice the side of this square and width three units less than the side of this square. If the two areas are equal what are the squares dimensions?
Answer by jorel1380(3719)

Let n be one side of the square. Then, the rectangle is 2n by n-3. So:
2n(n-3)=n(n)
2nē-6n-nē=0
nē=6n
n=6
The square is 6X6
☺☺☺☺